OpenStudy (anonymous):

When 2.0 A of current moves across a 1.0 MΩ resistor, then what is the drop in voltage (∆V) across the resistor? A. 2.0 × 10^0 V B. 2.0 × 10^2 V C. 2.0 × 10^3 V D. 2.0 × 10^6 V

OpenStudy (michele_laino):

here we have to apply the Ohm's law: \[\Delta V = RI = 2 \times 1 \times {10^6}\]

OpenStudy (michele_laino):

since 1MOhm= 10^6 Ohms

OpenStudy (anonymous):

ohh ok! so we get choice D as our solution?

OpenStudy (anonymous):

is that correct? :)

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay! thanks!

OpenStudy (michele_laino):

:)